Of 200cm2 of carbon(iv)oxide were collected over water at 18°c and 700mmHg, determine the volume of thr dry gas at STP


1
Expert's answer
2020-07-14T15:20:17-0400

The CO2 gas collected over water contains water vapors, which partial pressure is 16 mmHg at a given temperature. According to the partial pressure law:


"p(total)=p(H_{2}O)+p(CO_{2});"


"p(CO_{2})=p(total)-p(H_{2}O)=700mmHg-16mmHg=684mmHg(91193Pa);"


Thus,


"V_{STP}(CO_{2})=\frac{p(CO_{2})V(CO_{2})}{RT}*\frac{RT_{0}}{P_{0}}=\frac{91193Pa*2*10^{-4}m^{3}}{8.31\frac{J}{mol*K}*291K}*\frac{8.31\frac{J}{mol*K}*273K}{101325Pa}=1.69*10^{-4}m^{3}(169mL)."